Session: Nov 1997 (session 4)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (cattani at Fri Nov 7 11:12:59 MET 1997) Minor failures
Problems throughout the experiment: tape slides in forward passes Sometimes minutes of recording lost and some scan too. It was realized that the problem was the capstan: substituted with a new one.

Onsala (A. Polatidis at Fri Nov 7 11:42:30 MET 1997) Success
No problems known.

Torun (K.Borkowski at Fri Nov 7 12:59:15 MET 1997) Success

Noto (C.Trigilio at Fri Nov 7 13:01:21 MET 1997) Success
No problem during the observations.

Jodrell (Alastair Gunn at Sat Nov 8 17:43:06 MET 1997) Success
No known problems. Track 21 dead.

Effelsberg (dgraham at Mon Nov 10 09:06:37 MET 1997) Success

Calibration (Fredrik T Rantakyro at Tue Dec 9 15:03:42 MET 1997) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the NOV97 directory under the name vt906.antab Read the comments in the file to get further information.